# The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency

### v1

@@ Abstract
In econometrics, the Efficient Market Hypothesis posits that asset prices reflect all available information in the market. Several empirical investigations show that market efficiency drops when it undergoes extreme events. Many models for multivariate extremes focus on positive dependence, making them unsuitable for studying extremal dependence in financial markets where data often exhibit both positive and negative extremal dependence. To this end, we construct regular variation models on the entirety of $\mathbb{R}^{d}$ and develop a bivariate measure for asymmetry in the strength of extremal dependence between adjacent orthants. Our directional tail dependence (DTD) measure allows us to define the Efficient Tail Hypothesis (ETH)—an analogue of the Efficient Market Hypothesis—for the extremal behaviour of the market. Asymptotic results for estimators of DTD are described, and we discuss testing of the ETH via permutation-based methods and present novel tools for visualization. Empirical study of China’s futures market leads to a rejection of the ETH and we identify potential profitable investment opportunities. To promote the research of microstructure in China’s derivatives market, we open-source our high-frequency data, which are being collected continuously from multiple derivative exchanges.
@@ 1 Introduction
Market efficiency is a fundamental concept in econometrics, which indicates how fast the market can absorb new information and how swiftly market prices change to reflect the fair value of an asset. The Efficient Market Hypothesis, first introduced by Fama, (1970), states that, in an ideal situation, the market price of an asset reflects all available information in the market and must be equal to the asset’s fair value. In other words, when the market is fully efficient, the price returns are entirely unpredictable. The debate on the use of the Efficient Market Hypothesis for statistical analysis has been ongoing for decades, and the Efficient Market Hypothesis has been challenged by many empirical studies; see, e.g., Lo, (2004, 2005) and Malkiel, (2003). Various statistical methods have been proposed to test the Efficient Market Hypothesis (Choi,, 1999; Escanciano and Lobato,, 2009; Charles et al.,, 2011; Wang,, 2024). However, these methods are often based on the assumption that log-returns are normally distributed, while, in reality, the returns typically have much heavier tails and non-Gaussian dependencies, especially in a high-frequency setting (Cont,, 2001; De Domenico et al.,, 2023). Moreover, increasing evidence shows that the market is not always efficient, especially when it experiences extreme events during financial turmoil (Choi,, 2021); this kind of inefficiency in tails is often ignored by traditional Efficient Market Hypothesis tests.
@@
Many extremal dependence models have been developed within the last decade (see, e.g., Castro-Camilo et al.,, 2018; Huang et al.,, 2019; Engelke and Hitz,, 2020; Chan et al.,, 2022; Gong and Huser,, 2022; Gong et al.,, 2024), and such models are widely used in risk management and insurance. For example, Gong et al., (2024) study the extremal dependence between the depreciation of different currencies; Chan et al., (2022) examine the extremal dependence between trading volumes and returns in cryptomarkets; In the same vein, Van Oordt and Zhou, (2019) propose an estimator for linear coefficients of a bivariate regular variation model under extreme adverse market conditions; Bouaddi and Moutanabbir, (2022) study the asymmetry in contagion effects between the US stock market and other international markets by comparing extreme losses and gains; Lu et al., (2024) study the extreme co-movement between three decomposed oil prices by estimating a tail correlation coefficient. For a review of many more applications of extremal dependence models in financial applications, see Nolde and Zhou, (2021).
@@
Traditional extremal dependence models may not be adequate for studying market efficiency as they are often founded on an underlying assumption that data are regularly-varying in the positive orthant of $\mathbb{R}^{d}$ only, i.e., when all variables are positive and jointly heavy-tailed; for details on regular variation, see, e.g., Resnick, (2007), Chapter 6. Thus, such models can assess only positive extremal dependence, i.e., the dependence between variables as they jointly grow large. While joint losses may be studied by flipping the sign of the data, extremal dependence in all orthants (of $\mathbb{R}^{d}$) is important in the study of market efficiency, as we seek to model the joint occurrence of combinations of both losses and gains of different assets. Motivated by the lack of tools for testing market efficiency in extreme situations, we consider random vectors $\bm{X}:=(X_{1},\dots,X_{p})^{\top}\in\mathbb{R}^{p}$, that are regularly-varying on all orthants of $\mathbb{R}^{p}$. We derive mass balance constraints for these vectors to construct, without loss of generality, balanced regularly-varying random vectors in which both upper and lower tails are heavy-tailed and of unit scale. Furthermore, we define a new pairwise measure of directional tail dependence (DTD), denoted by $\lambda(\cdot,\cdot)$, which quantifies the strength of asymmetry in the extremal dependence exhibited in adjacent orthants of $\mathbb{R}^{d}$ for these balanced regularly-varying random vectors; this measure extends the popular pairwise Extremal Dependence Measure (EDM), proposed by Resnick, (2004). The terminology of asymmetric tail dependence has been broadly used in the literatures to describes the asymmetry between joint upper tails region and joint lower tails region (see, e.g., Echaust,, 2021; Gong and Huser,, 2022). To avoid confusion, we use the terminology of “directional (a)symmetry” here for describing the (a)symmetry of extremal dependence in two adjacent quadrants, as one variable plays a leading role in this framework for the occurrence of extreme events in the other variable. We propose two estimators for $\lambda(\cdot,\cdot)$ and study their asymptotic properties, as well as construct a non-parametric permutation test for the symmetric case where $\lambda(\cdot,\cdot)=0$. In financial markets, let $\bm{Y}\in\mathbb{R}^{p_{2}}$ represent the asset returns at a given time, and $\bm{X}\in\mathbb{R}^{p_{1}}$ represent the information that is measurable before the return of the assets. We then define the Efficient Tail Hypothesis (ETH), a new concept analogous to the Efficient Market Hypothesis, which corresponds to the situation where $\lambda(X_{i},Y_{j})=0$ for $i=1,\dots,p_{1},j=1,\dots,p_{2},$ where $p_{2}$ is the number of assets and $p_{1}$ is the number of indicators. We also provide a means of visualising directional tail asymmetry for high-dimensional random vectors by projecting values of $\lambda$ to the surface of a sphere.
@@
We apply our method to study the market-wide tail-efficiency of China’s futures market, using high-frequency data containing second-level market information that we mined for use in statistical analysis. China’s futures market has emerged as one of the most important markets globally, with transactions amounting to 80 trillion United States (U.S.) dollars annually and 70 types of commodity futures. An increasing body of research indicates that China’s derivatives market exerts a significant pricing influence on the global market (Arslanalp et al.,, 2016). Given its magnitude and influence, there is considerable potential for empirical studies of this market in order to validate theoretical results. However, despite China’s growing impact, there are limited accessible and user-friendly datasets available for researchers specializing in China’s derivatives market. While there are commercial data providers, such as Wind, Bloomberg, and Choice, their data are not open-source. To promote research on China’s derivatives market, we have gathered millisecond-level, high-frequency data since 2022 from multiple exchanges, totaling more than 1.3 Terabytes of data, and we provide it as open-source for academic purposes. A subset of these data have been considered previously by Jiang et al., (2022) in market-making research. In our analysis, we find that most of the time-leading tail dependencies are symmetric, i.e., $\lambda(X_{i},Y_{j})=0,$ but some inefficiencies persist. To capitalize on this finding, we construct an artificial dynamic portfolio consisting of assets that do not satisfy the conditions of the ETH and demonstrate that it can generate profits on out-of-sample data.
@@ 2 Multivariate extreme-value background
Definition 2.1 (Multivariate regular variation (§6, Resnick, (2007))).
@@
where $\oplus$ and $\circ$ are transformed-linear operators (see Cooley and Thibaud, (2019) for details) that satisfy $X_{1}\oplus X_{2}=t\{t^{-1}(X_{1})+t^{-1}(X_{2})\}$ and $a\circ X=t\{at^{-1}(X)\}$, for constant $a\in\mathbb{R}$ and ${t(\cdot)=\log\{1+\exp(\cdot)\}}$. The space $\mathcal{V}^{q}_{+}$ is a subspace of ${\rm RV}^{1}_{+}(2)$, and contains elements that can be spanned (using the transformed-linear operations) by a vector of independent regularly varying ${\rm RV}^{1}_{+}(2)$ random variables with non-negative coefficients. This space has two important properties, which make it conducive to studying the tails of random variables (see, e.g., the transformed linear model of Mhatre, (2022) and the partial tail correlation coefficients of Lee and Cooley, (2022) and Gong et al., (2024)): first, the corresponding angular measure $H_{\mathbf{X}}(\cdot)$ for any $\bm{X}\in{\rm RV}^{p}_{+}(2)$ can be approximated using a sequence of transformed-linear combinations of independent regularly varying random variables (Proposition 4; Cooley and Thibaud,, 2019) with increasing $q$. This implies that $\mathcal{V}^{q}_{+}$ is dense in ${\rm RV}^{p}_{+}(2)$ as $q\to\infty$. Second, the inner product in $\mathcal{V}^{q}_{+}$, i.e., $\langle X,Y\rangle:=\bm{a}_{x}^{T}\bm{a}_{y}$ where $\bm{a}_{x}$ and $\bm{a}_{y}$ are the coefficient vectors of $X$ and $Y$, respectively, in the transformed-linear space $\mathcal{V}^{q}_{+}$, can be related to the Extremal Dependence Measure (EDM). This measure was first introduced by Resnick, (2004) and later studied by Larsson and Resnick, (2012) and Cooley and Thibaud, (2019).
@@ 3 Methodology
In Section 3.1, we construct balanced regularly-varying random vectors on $\mathbb{R}^{p}$ which have both upper and lower tails of unit scale and allow extremal dependence to be modeled simultaneously for both tails. Such a setting appeals to fields where extremal events can happen in both directions, such as the financial sector. In Section 3.2, we discuss a method for visualizing directional asymmetric extremal dependence in such random vectors. In Section 3.3, we propose our measure of directional tail dependence. In Section 3.4, we introduce the Efficient Tail Hypothesis (ETH) and propose methods for testing it.
@@ 3.1 Balanced regular variation
A $p$-dimensional random vector $\bm{X}\in\mathbb{R}^{p}$ is regularly-varying with tail index $\alpha>0$, denoted by $\bm{X}\in{\rm RV}^{p}(\alpha)$, if $|\bm{X}|\in{\rm RV}_{+}^{p}(\alpha)$ with the normalizing sequence $b_{n}\to\infty$ and for all ${\bm{z}\in[0,\infty)^{p}\setminus\{\bm{0}\}}$ and $\bm{s}\in\{-1,1\}^{p}$,
@@
[…]
To focus on describing extremal dependence structure, we put the upper and lower tails on the same starting line and consider a “balanced” version of $\bm{X}\in{\rm RV}^{p}(\alpha)$, where the scale of the marginal upper and lower tails of $\bm{X}$ are all equal. Define scaling factors for upper and lower tails ${c^{+}_{i}:=\nu_{\bm{X}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\}\right)}$ and ${c^{-}_{i}:=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})}$ and consider $\lim_{n\rightarrow\infty}{\mathbb{P}\{X_{i}<-xb_{n}\}\over\mathbb{P}\{X_{i}>xb_{n}\}}={c^{-}_{i}\over c^{+}_{i}}$ for all $i=1,\dots,p$. It is possible that the tails are unbalanced in the sense that $c^{+}_{i}$ and $c^{-}_{i}$ are unequal for at least one of $i\in\{1,\dots,p\}$. To ensure that $c_{i}^{+}=c_{i}^{-}$ for all $i=1,\dots,p$, we use the transformation defined in Proposition 3.1 below and denote the resulting random vector as $\bm{X}\in{\rm BRV}^{p}(\alpha)$; see Definition 3.2.
@@
A $p$-dimensional regularly-varying random vector $\bm{X}\in{\rm RV}^{p}(\alpha)$ is balanced, denoted by $\bm{X}\in{\rm BRV}^{p}(\alpha)$, if and ${v_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\})}={v_{\bm{X}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\}\right)=1}$ for all $i=1\dots,p$.
@@
Recall that we consider $\alpha=2$ for constructing the inner product space. In practice, our consideration for data to be balanced regularly-varying is without loss of generality, as we can always find a marginal transformation of $\bm{X}\sim{\rm RV}^{p}(2)$ to ensure it satisfies the balance condition. One such transformation is the one in Proposition 3.1; another one is to standardize margins to symmetric Pareto with unit scale and shape 2; the corresponding distribution function is
$F(x)=\begin{cases}(\sqrt{2}-x)^{-2},&x\leq 0,\\ 1-(x+\sqrt{2})^{-2},&x>0.\end{cases}$ (3)
Whilst ${\rm BRV}^{p}(2)$ random vectors present a natural modeling framework for studying multivariate extremal dependence, it is not trivial to immediately quantify their associated extremal dependence. This is because the total mass of the angular measure $H_{\bm{X}}(\cdot)$ is not split evenly between the orthants of $\mathbb{R}^{p}$, and many of the standard tools available for multivariate extremes have been developed for ${\rm RV}^{p}_{+}(\alpha)$; see, e.g., Cooley and Thibaud, (2019); Lee and Cooley, (2021). Hence, we now consider a transformation of $\bm{X}\in{\rm BRV}^{p}(2)$ to the positive orthant which preserves information on the degree of asymmetry in the upper and lower tails. Such a representation is briefly mentioned in §6.5.5 of Resnick, (2007).
@@
This representation of $\bar{\bm{X}}$ has many advantages. First, the extremal dependence structure information in $\bm{X}$ is preserved in $\bar{\bm{X}}$, and any subvector of $\bar{\bm{X}}$ is still regularly-varying. Second, the total angular mass is altered in a deterministic way, which is helpful for proving theoretical properties of $\bar{\bm{X}}$. Third, the orthogonality property simplifies the analysis and visualization of extremal dependence (see Section 3.2).

### v2

# SReferences

@@ Abstract
In econometrics, the Efficient Market Hypothesis posits that asset prices reflect all available information in the market. Several empirical investigations show that market efficiency drops when it undergoes extreme events. Many models for multivariate extremes focus on positive dependence, making them unsuitable for studying extremal dependence in financial markets where data often exhibit both positive and negative extremal dependence. To this end, we construct regular variation models on the entirety of $\mathbb{R}^{d}$ and develop a bivariate measure for asymmetry in the strength of extremal dependence between adjacent orthants. Our directional tail dependence (DTD) measure allows us to define the Efficient Tail Hypothesis (ETH)—an analogue of the Efficient Market Hypothesis—for the extremal behaviour of the market. Asymptotic results for estimators of DTD are described, and we discuss testing of the ETH via permutation-based methods and present novel tools for visualization. An empirical study of China’s futures market leads to a rejection of the ETH and we identify potential profitable investment opportunities during extreme episodes. To promote the research of microstructure in China’s derivatives market, we open-source our high-frequency data, which are being collected continuously from multiple derivative exchanges.
@@ 1 Introduction
Market efficiency is a fundamental concept in econometrics, which indicates how fast the market can absorb new information and how swiftly market prices change to reflect the fair value of an asset. The Efficient Market Hypothesis, first introduced by Fama, (1970), states that, in an ideal situation, the market price of an asset reflects all available information in the market and must be equal to the asset’s fair value. In other words, when the market is fully efficient, the price returns are entirely unpredictable. The debate on the use of the Efficient Market Hypothesis for statistical analysis has been ongoing for decades, and the Efficient Market Hypothesis has been challenged by many empirical studies; see, e.g., Lo, (2004, 2005) and Malkiel, (2003). Various statistical methods have been proposed to test the Efficient Market Hypothesis (Lo and MacKinlay,, 1988; Choi,, 1999; Escanciano and Lobato,, 2009; Charles et al.,, 2011; Wang,, 2024). However, these methods are often based on the assumption that log-returns are normally distributed, while, in reality, the returns typically have much heavier tails and non-Gaussian dependencies, especially in a high-frequency setting (Cont,, 2001; De Domenico et al.,, 2023). Moreover, increasing evidence shows that the market is not always efficient, especially when it experiences extreme events during financial turmoil (Choi,, 2021); this kind of inefficiency in tails is often ignored by traditional Efficient Market Hypothesis tests.
@@
Many extremal dependence models have been developed within the last 20 years (see, e.g., Hartmann et al.,, 2004; Poon et al.,, 2004; Castro-Camilo et al.,, 2018; Huang et al.,, 2019; Engelke and Hitz,, 2020; Chan et al.,, 2022; Gong and Huser,, 2022; Gong et al.,, 2024), and such models are widely used in risk management and insurance. For example, Poon et al., (2004) study extremal dependence between returns of risky assets in major financial markets, and Hartmann et al., (2004) study cross-asset extremal dependence between stocks and government bonds and find that stock crises tend to be linked with booms in government bonds; this phenomenon is known as the “flight to quality”, and serves as an example of cross-directional extremal dependence. Other more recent studies include: Gong et al., (2024), who study extremal dependence between the depreciation of different currencies; Bouaddi and Moutanabbir, (2022), who study the asymmetry in contagion effects between the US stock market and other international markets by comparing extreme losses and gains; Chan et al., (2022), who examine the extremal dependence between trading volumes and returns in cryptomarkets; in the same vein, Van Oordt and Zhou, (2019) propose an estimator for linear coefficients of a bivariate regular variation model under extreme adverse market conditions; Lu et al., (2024) study the extreme co-movement between three decomposed oil prices by estimating a tail correlation coefficient. For a review of many more applications of extremal dependence models in finance, see Nolde and Zhou, (2021).
Empirical studies show that the efficiency of a market drops when it undergoes extreme events (Choi,, 2021), and extreme value theory is an effective framework to analyze extremal dependence between market assets (Poon et al.,, 2004). However, traditional extremal dependence models may not be adequate for studying market efficiency as they are often founded on an underlying assumption that data are regularly varying in the positive orthant of $\mathbb{R}^{d}$ only, i.e., when all variables are positive and jointly heavy-tailed; for details on regular variation, see, e.g., Resnick, (2007, Chapter 6). Thus, such models can assess only positive extremal dependence, i.e., the dependence between variables as they jointly grow large. While joint losses may be studied by flipping the sign of the data, extremal dependence in all orthants (of $\mathbb{R}^{d}$) is important in the study of market efficiency, as we seek to model the joint occurrence of combinations of both losses and gains of different assets. In this work, we develop a bivariate measure based on extreme value theory to quantify the difference in the strength of tail dependence in two adjacent quadrants; this allows us to investigate the direction of dependence during extremal events.
@@
We consider random vectors, $\bm{X}:=(X_{1},\dots,X_{p})^{\top}\in\mathbb{R}^{p}$, that are regularly-varying on all orthants of $\mathbb{R}^{p}$. Focusing on dependence modeling, we standardize the marginal distributions and derive mass balance constraints for these vectors to construct, without loss of generality (see Section 3.2), balanced regularly-varying random vectors in which both upper and lower tails are heavy-tailed and of unit scale. Furthermore, we define a new pairwise measure of directional tail dependence (DTD), denoted by $\lambda(\cdot,\cdot)$, which quantifies the strength of asymmetry in the extremal dependence exhibited in adjacent orthants of $\mathbb{R}^{d}$ for these balanced regularly-varying random vectors; this measure extends the popular pairwise Extremal Dependence Measure (EDM), proposed by Resnick, (2004). The terminology of “asymmetric” tail dependence has been broadly used in the literature to describe asymmetry between the joint upper-tail and joint lower-tail regions (see, e.g., Manner,, 2010; Jondeau,, 2016; Echaust,, 2021; Gong and Huser,, 2022). To avoid confusion, we use here the terminology of “directional (a)symmetry” for describing the (a)symmetry of extremal dependence in two adjacent quadrants, as one variable plays a leading role in this framework for the occurrence of extreme events in the other variable. We propose two estimators for $\lambda(\cdot,\cdot)$ and study their asymptotic properties, as well as construct a non-parametric permutation test for the symmetric case where $\lambda(\cdot,\cdot)=0$.
In financial markets, let $\bm{Y}\in\mathbb{R}^{p_{2}}$ represent the asset returns at a given time, and $\bm{X}\in\mathbb{R}^{p_{1}}$ represent the information that is measurable before the return of the assets. We then define the Efficient Tail Hypothesis (ETH), a new concept analogous to the Efficient Market Hypothesis, which corresponds to the situation where $\lambda(X_{i},Y_{j})=0$ for $i=1,\dots,p_{1},j=1,\dots,p_{2},$ where $p_{2}$ is the number of assets and $p_{1}$ is the number of indicators. We also provide a means of visualizing directional tail asymmetry for high-dimensional random vectors by projecting values of $\lambda$ to the surface of a sphere. From an economic standpoint, the ability to detect directional tail dependence, between extreme gains and losses, provides an informative view of the shock-propagation mechanism that occurs in financial markets during extreme periods. Moreover, a rejection of the ETH suggests the existence of assets that, when extreme, lead to extremes of other assets in the following time step; this can serve as an early-warning system for risk management or can be used to develop trading strategies. Therefore, our proposed methodology contributes to the literature on risk spillover analysis (see, e.g., Billio et al.,, 2012; Bollerslev et al.,, 2018; Jian et al.,, 2018) and lead-lag detection (see, e.g., Hou,, 2007; Tolikas,, 2018; Buccheri et al.,, 2021) by incorporating directional tail dependence.
We apply our method to study the market-wide tail-efficiency of China’s futures market, using high-frequency data containing second-level market information that we mined for use in statistical analyses. In our analysis, we find that most of the time-leading tail dependencies are symmetric, i.e., $\lambda(X_{i},Y_{j})=0,$ but some inefficiencies persist. To capitalize on this finding, we construct an artificial dynamic portfolio consisting of assets that do not satisfy the conditions of the ETH and demonstrate that it can generate profits using out-of-sample data.
@@
China’s futures market has emerged as one of the most important markets globally, with transactions amounting to 80 trillion United States (U.S.) dollars annually and 70 types of commodity futures. An increasing body of research indicates that China’s derivatives market exerts a significant pricing influence on the global market (Arslanalp et al.,, 2016). Given its magnitude and influence, there is considerable potential for empirical studies of this market in order to validate theoretical results. However, despite China’s growing impact, there are limited accessible and user-friendly datasets available for researchers specializing in China’s derivatives market. While there are commercial data providers, such as Wind, Bloomberg, and Choice, their data are not open-source. To promote research on China’s derivatives market, we have gathered millisecond-level, high-frequency data since 2022 from multiple exchanges, totaling more than 1.3 Terabytes of data, and we provide it as open-source for academic purposes. A subset of these data have been considered previously by Jiang et al., (2022) in market-making research.
@@ 2 Multivariate extreme-value background
Definition 2.1 (Multivariate regular variation (Resnick,, 2007; Chapter 6)).
@@
where $\oplus$ and $\circ$ are transformed-linear operators (see Cooley and Thibaud,, 2019, for details) that satisfy $X_{1}\oplus X_{2}=t\{t^{-1}(X_{1})+t^{-1}(X_{2})\}$ and $a\circ X=t\{at^{-1}(X)\}$, for constant $a\in\mathbb{R}$ and ${t(\cdot)=\log\{1+\exp(\cdot)\}}$. The space $\mathcal{V}^{q}_{+}$ is a subspace of ${\rm RV}^{1}_{+}(2)$, and contains elements that can be spanned (using the transformed-linear operations) by a vector of independent regularly varying ${\rm RV}^{1}_{+}(2)$ random variables with non-negative coefficients. This space has two important properties, which make it conducive to studying the tails of random variables (see, e.g., the transformed linear model of Mhatre, (2022) and the partial tail correlation coefficients of Lee and Cooley, (2022) and Gong et al., (2024)): first, the corresponding angular measure $H_{\mathbf{X}}(\cdot)$ for any $\bm{X}\in{\rm RV}^{p}_{+}(2)$ can be approximated using a sequence of transformed-linear combinations of independent regularly varying random variables (Proposition 4; Cooley and Thibaud,, 2019) with increasing $q$. This implies that $\mathcal{V}^{q}_{+}$ is dense in ${\rm RV}^{p}_{+}(2)$ as $q\to\infty$. Second, the inner product in $\mathcal{V}^{q}_{+}$, i.e., $\langle X,Y\rangle:=\bm{a}_{x}^{T}\bm{a}_{y}$ where $\bm{a}_{x}$ and $\bm{a}_{y}$ are the coefficient vectors of $X$ and $Y$, respectively, in the transformed-linear space $\mathcal{V}^{q}_{+}$, can be related to the Extremal Dependence Measure (EDM). This measure was first introduced by Resnick, (2004) and later studied by Larsson and Resnick, (2012) and Cooley and Thibaud, (2019).
@@ 3 Methodology
In Section 3.1, we construct balanced regularly varying random vectors on $\mathbb{R}^{p}$ which have both upper and lower tails of unit scale and allow extremal dependence to be modeled simultaneously for both tails. Such a setting appeals to fields where extremal events can happen in both directions, such as the financial sector. In Section 3.2, we demonstrate various marginal transformations that can be applied to data to produce balanced regularly varying random variables. In Section 3.3, we discuss a method for visualizing directional asymmetric extremal dependence in such random vectors. In Section 3.4, we propose our measure of directional tail dependence. In Section 3.5, we introduce the Efficient Tail Hypothesis (ETH) and propose methods for testing it.
@@ 3.1 Balanced regular variation
A $p$-dimensional random vector $\bm{X}\in\mathbb{R}^{p}$ is regularly varying with tail index $\alpha>0$, denoted by $\bm{X}\in{\rm RV}^{p}(\alpha)$, if $|\bm{X}|\in{\rm RV}_{+}^{p}(\alpha)$ with the normalizing sequence $b_{n}\to\infty$ and for all ${\bm{z}\in[0,\infty)^{p}\setminus\{\bm{0}\}}$ and $\bm{s}\in\{-1,1\}^{p}$,
@@
The properties of $\bm{X}\in{\rm RV}^{p}(\alpha)$ similarly involve a radial-angular decomposition, as previously described in Section 2 for ${\rm RV}^{p}_{+}(\alpha)$ random variables. The difference here is that the limit and angular mass measures are now defined on the punctured real space, $\mathbb{E}^{p}:=\mathbb{R}^{p}\setminus\{\bm{0}\},$ and the full unit $(p-1)$-sphere, respectively. Similarly (to ${\rm RV}^{p}_{+}(\alpha)$ random variables), the angular mass measure can be normalized to a unique probability measure $N_{\bm{X}}(\cdot):=H_{\bm{X}}(\cdot)/m$, with the normalizing constant $m=H_{\bm{X}}(\mathbb{S}^{p-1})$ representing the total angular mass on the unit sphere.
To focus on describing extremal dependence structure, we apply a marginal scaling of $\bm{X}$ to ensure that the scale of all of its marginal upper and lower tails are equal, and consider a “balanced” version of $\bm{X}\in{\rm RV}^{p}(\alpha)$. We define scaling factors for upper and lower tails by ${c^{+}_{i}:=\nu_{\bm{X}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\}\right)}$ and ${c^{-}_{i}:=\nu_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\})},$ respectively, and consider $\lim_{n\rightarrow\infty}{\mathbb{P}\{X_{i}<-xb_{n}\}\over\mathbb{P}\{X_{i}>xb_{n}\}}={c^{-}_{i}\over c^{+}_{i}}$ for all $i=1,\dots,p$. It is possible that the tails are unbalanced in the sense that $c^{+}_{i}$ and $c^{-}_{i}$ are unequal for at least one of $i\in\{1,\dots,p\}$. To ensure that $c_{i}^{+}=c_{i}^{-}$ for all $i=1,\dots,p$, we use the transformation defined in Proposition 3.1 below and denote the resulting random vector as $\bm{X}\in{\rm BRV}^{p}(\alpha)$; see Definition 3.2.
@@
A $p$-dimensional regularly varying random vector $\bm{X}\in{\rm RV}^{p}(\alpha)$ is balanced, denoted by $\bm{X}\in{\rm BRV}^{p}(\alpha)$, if ${v_{\bm{X}}(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\})}={v_{\bm{X}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\}\right)=1}$ for all $i=1\dots,p$.
@@
Recall that we consider $\alpha=2$ for constructing the inner product space. Whilst ${\rm BRV}^{p}(2)$ random vectors present a natural modeling framework for studying multivariate extremal dependence, it is not trivial to immediately quantify their associated extremal dependence. This is because the total mass of the angular measure $H_{\bm{X}}(\cdot)$ is not split evenly between the orthants of $\mathbb{R}^{p}$, and many of the standard tools available for multivariate extremes have been developed for ${\rm RV}^{p}_{+}(\alpha)$; see, e.g., Cooley and Thibaud, (2019); Lee and Cooley, (2021). Hence, we now consider a transformation of $\bm{X}\in{\rm BRV}^{p}(2)$ to the positive orthant which preserves information on the degree of asymmetry in the upper and lower tails. Such a representation is briefly mentioned in §6.5.5 of Resnick, (2007).
@@
This representation of $\bar{\bm{X}}$ has many advantages. First, the extremal dependence structure information in $\bm{X}$ is preserved in $\bar{\bm{X}}$, and any subvector of $\bar{\bm{X}}$ is still regularly varying. Second, the total angular mass is altered in a deterministic way, which is helpful for proving theoretical properties of $\bar{\bm{X}}$. Third, the orthogonality property simplifies the analysis and visualization of extremal dependence (see Section 3.3).
3.2 Marginal transformation to balanced regular variation
In practice, the upper and lower tails of a random variable may exhibit different tail heaviness, and the margins of our asset return data, denoted here by $\mathbf{R}\in\mathbb{R}^{p},$ may not be balanced, i.e., they may have different scales or tail-heaviness. However, the methodology in this paper remains applicable after standardizing the data margins to ensure balanced regular variation is satisfied. The pre-standardization of data to common margins is common practice for modeling of multivariate extremes (see discussion by, e.g., Naveau and Segers,, 2024). We describe here two marginal transformations of $\mathbf{R}$ to a balanced regularly varying random vector ${\rm BRV}^{p}(2)$: the first assumes that the upper and lower tails of all components of $\mathbf{R}$ are marginally regularly varying (we find this to be the case for the data in our application; see Section 5 and Section D.2 of the Supplementary Material), whilst the second is applicable for general data margins.
The first transformation follows from Proposition 3.1: let $\lim_{r\rightarrow\infty}\mathbb{P}\{R_{i}>r\}/r^{-\alpha_{i}^{+}}=c_{i}^{+}$ and $\lim_{r\rightarrow\infty}\mathbb{P}\{R_{i}<-r\}/r^{-\alpha_{i}^{-}}=c_{i}^{-}$, where $c^{+}_{i},c^{-i}_{i},\alpha^{+}_{i},\alpha^{-}_{i}>0$ for $i=1,\dots,p$, i.e., the upper and lower tails of each component of $\mathbf{R}$ are regularly varying, albeit with different tail index and scale. We can transform $\mathbf{R}$ to $\bm{X}$ satisfying the balanced regular variation condition, by setting
$X_{i}=\begin{cases}(c_{i}^{+})^{-\frac{1}{2}}\times R_{i}^{{\alpha}_{i}^{+}/2},\quad&\text{for }R_{i}\geq 0,\\ -(c_{i}^{-})^{-\frac{1}{2}}\times|R_{i}|^{{\alpha}_{i}^{-}/2},\quad&\text{for }R_{i}<0,\end{cases}$ (3)
for all $i=1,\dots,p$. In practice, we need to estimate the tail indices, $\alpha_{i}^{+}$ and $\alpha_{i}^{-}$, e.g., using the Hill, (1975) estimator, and the scaling coefficients, $c_{i}^{+}$ and $c_{i}^{-}$, using $q(r^{(1-q)}_{i})^{\alpha_{i}^{+}}$ and $-q(r^{(q)}_{i})^{\alpha_{i}^{-}},$ respectively, where $r^{(q)}_{i}$ denotes the $q$-quantile of $R_{i}$ for $q>0$ close to zero.
If all data margins cannot reasonably be assumed to be regularly varying, one can instead apply an empirical rank-based transformation of the data onto uniform margins before transforming them back to symmetric Pareto with unit scale and shape 2; this has distribution and quantile functions given by
$F(x)$ =$=\begin{cases}\displaystyle(\sqrt{2}-x)^{-2},&x\leq 0,\\[10.0pt] \displaystyle 1-(x+\sqrt{2})^{-2},&x>0,\end{cases}\quad F^{-1}(q)$ =$=\begin{cases}\displaystyle\sqrt{2}-q^{-1/2},&0<q\leq\tfrac{1}{2},\\[10.0pt] \displaystyle(1-q)^{-1/2}-\sqrt{2},&\tfrac{1}{2}<q<1,\end{cases}$ (4)
respectively. Note that the distribution function (and its corresponding inverse) in (4) is asymptotically equivalent to the distribution (inverse) function of $X_{i}$ in (3).
The parametric tail-index-based transformation in Equation (3) requires that the original data have regularly varying tails, whilst the rank-based transformation in Equation (4) has no such requirement (and is suitable for general data margins); however, it might lack efficiency when estimating the tail behavior due to the scarcity of extreme events. Sensitivity of the dependence analysis to the choice of marginal transformation should be investigated; see Section 5.
3.3 Visualizing directional tail asymmetry